# arXiv:2602.24153v1[math.AG] 27 Feb 2026

### v1

@@ Abstract.
Let $f$ be a defining polynomial of a reduced projective curve $C\subset{\bf P}^{2}$ having only weighted homogeneous singularities. We show that the strong monodromy conjecture for $f$ follows rather easily from arxiv:1609.0480v11 using a formula of Denef and Loeser for the Newton-nondegenerate case with three variables (which can be deduced in the applied case from the one for the two variable case) together with known results about the strong monodromy conjecture in the two variable case. Here an amazing cancellation occurs so that possible counterexamples fail.
@@ Introduction
In the case d= 3,4d\,{=}\,3,4, the existence of such a vector field implies that either the conclusion of Theorem 2 is satisfied or the curve $C$ is a union of a line and a conic or two conics both intersecting at one point so that 3/d-3/d is a root of the Bernstein-Sato polynomial at a singular point of $C$, see [Sa 25, Remark 8.5]. It is then easy to reduce the strong monodromy conjecture  for the local topological zeta function  $Z^{\rm top}_{f,0}(s)$ (claiming that a pole of $Z^{\rm top}_{f,0}(s)$ is a root of the Bernstein-Sato polynomial $b_{f}(s)$, see [DeLo 92], [Lo 88], [BaVe 26]) to the following (by using known results in the two variable case).
@@
Note finally that the strong monodromy conjecture has been shown quite recently in the not necessarily reduced curve case under the hypothesis that the associated reduced curve has only weighted homogeneous singularities, see [BaVe 26].
@@ 1. Topological zeta function
Let $\pi\,{:}\,\widetilde{Y}\,{\to}\,Y\,{:=}\,{\mathbb{C}}^{n}$ be an embedded resolution of singularities of $X\,{:=}\,f^{-1}(0)$ with $D_{i}$ the irreducible components of $\widetilde{X}\,{:=}\,\pi^{-1}(X)$. Let $N_{i}$, $\nu_{i}\thinspace{-}\thinspace 1$ be the multiplicities of $\pi^{*}f$ and $\pi^{*}{\rm d}x_{1}{\wedge}\cdots{\wedge}{\rm d}x_{n}$ respectively. Set $D_{I}\,{:=}\,\hbox{$\bigcap$}_{i\in I}\,D_{i}$, $D^{\circ}_{I}\,{:=}\,D_{I}\,{\setminus}\,\hbox{$\bigcup$}_{i\notin I}R_{i}$. The local topological zeta function is defined by
@@
Im the curve case with n= 2n\,{=}\,2, the Euler number $\chi\bigl(D^{\circ}_{I}\cap\pi^{-1}(0)\bigr)$ vanishes if |I|= 1|I|\,{=}\,1 and $|\hbox{$\bigcup$}_{j\neq i}D_{i}\cap D_{j}|\,{=}\,2$, but the summands $\hbox{$\prod$}_{i\in I}\,(N_{i}s\thinspace{+}\thinspace\nu_{i})^{-1}$ for |I|= 2|I|\,{=}\,2 remain if $\hbox{$\bigcap$}_{i\in I}\,D_{i}\neq\emptyset$ and they may cancel each other out by using
@@ 2. Newton-nondegenerate case
We explain the notation only in the case n= 3n\,{=}\,3. (This is similar and easier when n= 2n\,{=}\,2.) Firstly $d_{\sigma}!\thinspace{\rm Vol}(\sigma)$ is the normalized volume  of $\sigma$. In the case $d_{\sigma}\,{=}\,0$ or 1, it is equal to 1 or $|\sigma\cap{\mathbb{N}}^{3}|\thinspace{-}\thinspace 1$ respectively. If $d_{\sigma}\,{=}\,2$ and $\sigma$ is simplicial so that the vertices of $\sigma$ are three points $a_{1},a_{2},a_{3}\,{\in}\,{\mathbb{N}}^{3}$, it coincides with the absolute value of the determinant of the matrix $(a_{1},a_{2},a_{3})$ divided by the lattice-theoretic distance  $N_{\sigma}$ between $\sigma$ and 0. The latter is defined by $N_{\sigma}\,{:=}\,\ell_{\sigma}(a_{i})$ for any $i\,{\in}\,\{1,2,3\}$ with $\ell_{\sigma}$ the linear function with integral coefficients whose greatest common divisor is 1 and such that $\ell_{\sigma}$ is orthogonal to $\sigma$, that is, $\sigma\,{\subset}\,\ell_{\sigma}^{-1}(N_{\sigma})$. We put $\nu_{\sigma}\,{:=}\,\ell_{\sigma}(1,1,1)$, which is equal to the sum of the coefficients of $\ell_{\sigma}$. In general the normalized volume is defined by taking a partition of $\sigma$. Note that $\ell_{\sigma}$, $N_{\sigma}$, $\nu_{\sigma}$ are well defined for non-simplicial 2-dimensional $\sigma$. These can be defined also for non-compact faces (after taking the projection to ${\mathbb{R}}^{2}$ or $\mathbb{R}$} if necessary), where $N_{\sigma}\,{=}\,0$ if two of the coefficients of $\ell_{\sigma}$ vanishes.
@@
This can be defined more generally by choosing a simplicial partition of the cone spanned by $\ell_{\sigma_{1}},\dots,\ell_{\sigma_{r}}$. It is quite amazing that the sum is independent of the choice of a partition, see [DeLo 92, Lemma 5.1.1].
@@ 3. Proof of Theorem 3
Let $P,Q$ be the vertices of the unique compact face $\sigma_{0}$ of $\Gamma_{\!+}(f)$, which is 1-dimensional. Dividing $f$ by a monomial of highest degree which divides $f$, we see that their coordinates are given by
@@
There are five 2-dimensional faces $\sigma_{i}$ of $\Gamma_{\!+}(f)$ with $i\,{\in}\,[1,5]$. The coefficients of their corresponding linear functions $\ell_{i}$ for $i\,{\in}\,[1,5]$ are respectively
@@
$\quad\hbox{with}\quad m\,{:=}\,ci\thinspace{+}\thinspace ak\thinspace{+}\thinspace acr,\quad n\,{:=}\,cj\thinspace{+}\thinspace bk\thinspace{+}\thinspace bcr.$
@@
Indeed, for the first two equalities it is enough to calculate the minor determinants of $\bigl(\begin{smallmatrix}c&0&a\\ 0&c&b\end{smallmatrix}\bigr)$ and the determinant of $\bigl(\begin{smallmatrix}0&0&1\\ c&0&a\\ 0&c&b\end{smallmatrix}\bigr)$. For the last one, we verify that the 2-dimensional faces containing $Q$ are $\sigma_{1}$, $\sigma_{2}$, $\sigma_{5}$, $\sigma_{4}$, and compute the determinants of $\bigl(\begin{smallmatrix}1&0&0\\ 0&1&0\\ 0&c&b\end{smallmatrix}\bigr)$ and $\bigl(\begin{smallmatrix}0&c&b\\ c&0&a\\ 1&0&0\end{smallmatrix}\bigr)$. (There is another expression, but they coincide according to a computer algebra program.)
@@
This can be used to prove the formula (2.1) for our case by comparing it with the one for the two variable case. What is very amazing is that the numerator  of the “global” topological zeta functions of $C$ is divisible  by $s\thinspace{+}\thinspace 3/d$ so that the pole 3/d-3/d disappears. Note that we need the division by $ds\thinspace{+}\thinspace 3$ for the passage from the “global” zeta functions of $C$ to $Z^{\rm top}_{f,0}(s)$ (via the blowup at 0) by definition. Using Macaulay2 for instance, this can be seen as follows.
@@
Here $Z,Z1,Z2$ are local topological zeta functions of $f$ and the two singular points of $C$, and $R$ is the “remaining” part of the “global” zeta function of $C$. This shows that $Z1\thinspace{+}\thinspace Z2\thinspace{+}\thinspace R$ is divisible by $ds\thinspace{+}\thinspace 3$, and the quotient coincides with $Z$ in the fraction field of the polynomial ring. This gives also a “proof” of (2.1) for our case assuming it in the two variable case (and using a computer).

### v2

# arXiv:2602.24153v2[math.AG] 02 Mar 2026

@@ Abstract.
Let $f$ be a defining polynomial of a reduced projective curve $C\subset{\bf P}^{2}$ having only weighted homogeneous singularities. We show that the strong monodromy conjecture for $f$ follows rather easily from arxiv:1609.04801v11 using a formula of Denef and Loeser for the Newton-nondegenerate case with three variables (which can be deduced in the applied case from the one for the two variable case) together with known results about the strong monodromy conjecture in the two variable case. Here an amazing cancellation occurs so that possible counterexamples fail.
@@ Introduction
In the case d= 3,4d\,{=}\,3,4, the existence of such a vector field implies that either the conclusion of Theorem 2 is satisfied or the curve $C$ is a union of a line and a conic or two conics both intersecting at one point so that 3/d-3/d is a root of the Bernstein-Sato polynomial at a singular point of $C$, see [Sa 25, Remark 8.5]. It is then easy to reduce the strong monodromy conjecture  for the local topological zeta function  $Z^{\rm top}_{f,0}(s)$ (claiming that a pole of $Z^{\rm top}_{f,0}(s)$ is a root of the Bernstein-Sato polynomial $b_{f}(s)$, see [DeLo 92], [BaVe 26]) to the following (by using known results in the two variable case [Lo 88], see Remark 3.1 below).
@@
Note finally that the strong monodromy conjecture has been shown quite recently in the not necessarily reduced curve case under the hypothesis that the associated reduced curve has only weighted homogeneous singularities, see [BaVe 26], where the vanishing of the residue at 3/d-3/d is proved in the needed case. Unfortunately it does not seem to be mentioned there that the following assertion actually holds.
Proposition 1.
Let $Z\,{\subset}\,{\mathbb{P}}^{n-1}$ be a reduced projective hypersurface defined by a polynomial $f$ with $n\,{\geqslant}\,4$. If the hypersurface $Z$ has only homogeneous isolated singularities which are not ordinary double points, then the strong monodromy conjecture holds for $f$.
This can be reduced to [BaVe 26, Theorem C] using Proposition 3.3 below, since the strong monodromy conjecture for homogeneous polynomials with isolated singularities is easy. This proposition seems to hold also in the non-reduced case.
@@ 1. Topological zeta function
Let $\pi\,{:}\,\widetilde{Y}\,{\to}\,Y\,{:=}\,{\mathbb{C}}^{n}$ be an embedded resolution of singularities of $D\,{:=}\,f^{-1}(0)$ with $D_{i}$ the irreducible components of $\widetilde{D}\,{:=}\,\pi^{-1}(D)$. Let $N_{i}$, $\nu_{i}\thinspace{-}\thinspace 1$ be the multiplicities of $\pi^{*}f$ and $\pi^{*}{\rm d}x_{1}{\wedge}\cdots{\wedge}{\rm d}x_{n}$ respectively. Set $D_{I}\,{:=}\,\hbox{$\bigcap$}_{i\in I}\,D_{i}$, $D^{\circ}_{I}\,{:=}\,D_{I}\,{\setminus}\,\hbox{$\bigcup$}_{i\notin I}R_{i}$. The local topological zeta function is defined by
@@
In the curve case with n= 2n\,{=}\,2, the Euler number $\chi\bigl(D^{\circ}_{I}\cap\pi^{-1}(0)\bigr)$ vanishes if |I|= 1|I|\,{=}\,1 and $|\hbox{$\bigcup$}_{j\neq i}D_{i}\cap D_{j}|\,{=}\,2$, but the summands $\hbox{$\prod$}_{i\in I}\,(N_{i}s\thinspace{+}\thinspace\nu_{i})^{-1}$ for |I|= 2|I|\,{=}\,2 remain if $\hbox{$\bigcap$}_{i\in I}\,D_{i}\neq\emptyset$ and they may cancel each other out by using
@@ 2. Newton-nondegenerate case
We explain the notation only in the case n= 3n\,{=}\,3, see [DeLo 92], [LeVP 11] for the general case. (This is similar and easier when n= 2n\,{=}\,2.) Firstly $d_{\sigma}!\thinspace{\rm Vol}(\sigma)$ is the normalized volume  of $\sigma$. In the case $d_{\sigma}\,{=}\,0$ or 1, it is equal to 1 or $|\sigma\cap{\mathbb{N}}^{3}|\thinspace{-}\thinspace 1$ respectively. If $d_{\sigma}\,{=}\,2$ and $\sigma$ is simplicial so that the vertices of $\sigma$ are three points $a_{1},a_{2},a_{3}\,{\in}\,{\mathbb{N}}^{3}$, it coincides with the absolute value of the determinant of the matrix $(a_{1},a_{2},a_{3})$ divided by the lattice-theoretic distance  $N_{\sigma}$ between $\sigma$ and 0. The latter is defined by $N_{\sigma}\,{:=}\,\ell_{\sigma}(a_{i})$ for $i\,{\in}\,\{1,2,3\}$ with $\ell_{\sigma}$ the linear function with integral coefficients whose greatest common divisor is 1 and such that $\ell_{\sigma}$ is orthogonal to $\sigma$, that is, $\sigma\,{\subset}\,\ell_{\sigma}^{-1}(N_{\sigma})$. We put $\nu_{\sigma}\,{:=}\,\ell_{\sigma}(1,1,1)$, which is equal to the sum of the coefficients of $\ell_{\sigma}$. In general the normalized volume is defined by taking a partition of $\sigma$. Note that $\ell_{\sigma}$, $N_{\sigma}$, $\nu_{\sigma}$ are well defined for non-simplicial 2-dimensional $\sigma$. These can be defined also for non-compact faces (after taking the projection to ${\mathbb{R}}^{2}$ or $\mathbb{R}$} if necessary), where $N_{\sigma}\,{=}\,0$ or $1$ if two of the coefficients of $\ell_{\sigma}$ vanishes (since $f$ is reduced).
@@
This can be defined more generally by choosing a simplicial partition of the cone spanned by $\ell_{\sigma_{1}},\dots,\ell_{\sigma_{r}}$. It is quite amazing that the sum is independent of the choice of a partition, see [DeLo 92, Lemma 5.1.1]. (This can be verified by using a computer in the situation of the proof of Theorem 3, see Remark 3.2 below.)
@@ 3. Proof of Theorem 3
Let $P,Q$ be the vertices of the unique compact face $\sigma_{0}$ of $\Gamma_{\!+}(f)$, which is 1-dimensional. Taking a monomial $x^{i}y^{j}z^{k}$ of highest degree which divides $f$, we see that their coordinates are given by
@@
There are five 2-dimensional faces $\sigma_{i}$ of $\Gamma_{\!+}(f)$ with $i\,{\in}\,[1,5]$. The coefficients of their corresponding linear functions $\ell^{(i)}$ for $i\,{\in}\,[1,5]$ are respectively
@@
$\hbox{with}\quad\quad m\,{:=}\,ci\thinspace{+}\thinspace ak\thinspace{+}\thinspace acr,\quad n\,{:=}\,cj\thinspace{+}\thinspace bk\thinspace{+}\thinspace bcr.$
@@
Indeed, for the first two equalities it is enough to calculate the minor determinants of $\bigl(\begin{smallmatrix}c&0&a\\ 0&c&b\end{smallmatrix}\bigr)$ and the determinant of $\bigl(\begin{smallmatrix}0&0&1\\ c&0&a\\ 0&c&b\end{smallmatrix}\bigr)$. For the last one, we verify that the 2-dimensional faces containing $Q$ are $\sigma_{1}$, $\sigma_{2}$, $\sigma_{5}$, $\sigma_{4}$, and compute the determinants of $\bigl(\begin{smallmatrix}1&0&0\\ 0&1&0\\ 0&c&b\end{smallmatrix}\bigr)$ and $\bigl(\begin{smallmatrix}0&c&b\\ c&0&a\\ 1&0&0\end{smallmatrix}\bigr)$. (Recall that there is another expression, but these coincide according to a computer algebra program, see Remark 3.2 below.)
@@
Remark 3.2.
The above argument can be used to prove the formula (2.1) for our case by comparing it with the one for the two variable case. What is very amazing is that the numerator  of the (modified) “global” topological zeta functions of $C$ (which contains the contribution from ${\mathbb{P}}^{2}\,{\setminus}\,C$) is divisible  by $s+3/d$ so that the pole 3/d-3/d disappears. Note that we need the division by $ds\thinspace{+}\thinspace 3$ for the passage from the (modified) “global” zeta functions of $C$ to $Z^{\rm top}_{f,0}(s)$ (via the blowup at 0) by definition. Using Macaulay2 for instance, this can be seen as follows.
@@
Here Z, Z1, Z2 are the local topological zeta functions of $f$ and the two singular points of $C$, and R is the “remaining” part of the (modified) “global” zeta function of $C$. (Notice that R is independent of $k\,{\in}\,\{0,1\}$, since $\{z\,{=}\,0\}\,{\setminus}\,{\rm Sing}\,C\cong{\mathbb{C}}^{*}$. In the cases (i,j)=(0,0)(i,j)\,{=}\,(0,0), $(1,0)$, and $(1,1)$, R is the contribution from ${\mathbb{P}}^{2}\,{\setminus}\,C$, $\{x\,{=}\,0\}\,{\setminus}\,{\rm Sing}\,C\,({\cong}\,{\mathbb{C}})$, and [0: 0: 1][0\,{:}\,0\,{:}\,1] respectively.) This computation shows that Z1+Z2+R is divisible by d*s+3, and the quotient coincides with Z in the field of fractions of the polynomial ring. This gives also a “proof” of (2.1) for our case assuming it in the two variable case (and using a computer algebra program).
It is not recommended to remove ; at the end of the definition of Z. If one prefers, S=QQ[a,b,i,j,k,r,s]; may be replaced by S=QQ[s]; a=2; b=3; i=0; j=0; k=0; r=1; for instance. The coincidence with another expression of $J_{Q}(s)$ can be seen by the following:
b/(i*s+1)/(j*s+1)/(n*s+b+c) + a*c/(n*s+b+c)/(m*s+a+c)/(i*s+1) -
(b*c/(j*s+1)/(n*s+b+c)/(m*s+a+c) + a/(m*s+a+c)/(i*s+1)/(j*s+1))
We note here the following (which is needed for a proof of Proposition 1).
Proposition 3.3.
Assume a reduced projective hypersurface $Z$ of dimension at least $2$ and degree at least $3$ has only homogeneous isolated singularities which are not ordinary double points, and there is a nonzero linear function $\ell(\nu_{1},\dots,\nu_{n})\,{=}\,\hbox{$\sum$}_{i=1}^{n}\,c_{i}\nu_{i}$ on ${\mathbb{R}}^{n}$ with rational coefficients $c_{i}$ such that ${\rm Supp}_{\bf x}\thinspace f\,{\subset}\,\ell^{-1}(0)$ with $f$ a defining polynomial of $Z$. Then $Z$ is a cone.
Proof.
[…]
We thus get that the $c_{j}$ are independent of $j\,{\neq}\,i$. These must vanish, since $f^{(j)}$ has non-isolated singularities otherwise. Then only $c_{i}$ is nonzero, that is, $f$ is a polynomial of $x_{j}$ for $j\,{\neq}\,i$. This contradicts the assumption that $Z$ is not a cone. So the assertion follows. ∎